# Sublogarithmic Distributed Vertex Coloring with Optimal Number of Colors

### 30 Mar 2026 v1

@@ 2 Technical Overview
2.1 Color Coverage (CC) & $\Pi$-ous subgraphs
@@
To circumvent this issue, Molloy and Reed constrain all earlier steps of the coloring algorithm so that the coloring outside each clique behaves sufficiently randomly. Concretely, this means that for every clique and every color, only a small number of vertices are forbidden from using that color. We formalize this requirement as the color coverage (henceforth CC) property. Ensuring that CC holds until coloring the dense parts of the graph is essential to completing the coloring. See Section 4.2 for the precise definition of CC.
@@
The CC condition can be maintained if all earlier coloring steps are sufficiently random. If each dense node has at most $U$ neighbors outside its clique, and each of them picks one random color out of $U\cdot\Delta^{0.22}$ colors, then few of these outside neighbors pick the same color, thereby maintaining the CC condition. Thus, the key to ensuring the CC condition in each step of the algorithm is to maintain large lists of available colors $L(v)$ (colors that do not currently appear in the neighborhood of $v$) compared to the number of outside neighbors of the cliques. We capture this condition through $\Pi$, which we state in a simplified form as follows:
@@
Subgraphs satisfying $\Pi$ are called $\Pi$-ous (pronounced pious). We can color a $\Pi$-ous subgraph $H$ while maintaining CC by a combination of iterated random color trials (Rct) and multi color trials (Mct). Condition (b) of $\Pi$ ensures that the list of available colors for each vertex of $H$ has $U\cdot\Delta^{0.22}$ slack that remains and ensures a sufficiently large list regardless of how the coloring of $H$ proceeds. Slack does not decrease because coloring a neighbor removes both one competing vertex and one available color, leaving their difference—the slack—unchanged or increased.
@@ 2.2 Overview of the Coloring Algorithm
Our algorithm centers on identifying $\Pi$-ous subgraphs that can be colored while maintaining CC, thereby ensuring that the dense vertices can be colored last. We build on the structural decomposition of Molloy and Reed, which nearly [^4] The subgraphs induced by $B_{H}$ and $B_{L}$ are $\Pi$-ous by construction, however the subgraph induced by $S$ is not $\Pi$-ous right away; see Section 6 for more details. partitions the sparser regions of the graph into $\Pi$-ous subgraphs [45]. More precisely, they reduced the problem of finding a $c$-coloring of $G$ to the problem of finding a $c$-coloring of a graph $F$ with a very specific structure. Bamas and Esperet also showed that $F$ can be constructed from $G$ in $O(1)$ rounds in the $\mathsf{LOCAL}$ model, and that one can efficiently recover a $c$-coloring of $G$ from a $c$-coloring of $F$ in the distributed setting [4].
@@
To achieve a sublogarithmic distributed runtime, we color cliques in $A_{H}$ and $A_{L}$ respectively in parallel. For each clique $A_{i}$, we assign a random permutation of the colors to its vertices avoiding any conflicts inside the clique. With this process, only $O(\sqrt{\Delta})$ vertices per clique conflict with external neighbors. Each such unhappy vertex $v$ seeks a partner $u$ in $A_{i}$ so that both become happy if they swap their colors. CC crucially guarantees a large supply of valid swap candidates.
@@ 2.3 Smaller Degrees and the Lovász Local Lemma
As one randomly colors vertices in the pre-shattering phase, one may be forced to uncolor certain vertices. For instance, if a color appears too frequently around a clique and CC is violated. This may interfere with the progress measures of our algorithm, which are typically formulated as additional bad events. For instance, uncoloring vertices may violate an event ensuring a constant-factor decrease in the uncolored degree. Such events are usually resolved during the post-shattering phase.
@@
Since we only pick candidates in pre- and post-shattering using separate budgets, this increases the number of neighbors with a given color by a factor of at most two compared to having one single budget for both steps. Conveniently, the same analysis bounds the probability of breaking CC for the pre- and post-shattering.
@@ 3. III:
By (S2), (S4), and the union bound over all sets $A$ with $v$ as a vertex, we have that
@@
By (S3), whether $A\in\mathcal{A}^{\prime}$ depend only on the random variables $X_{u}$ of vertices in $A$, which has diameter one in $H$. Whether $v\in N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}}(\mathcal{A}^{\prime})$ depends on whether $A\in\mathcal{A}^{\prime}$ for all the $A$ with a vertex in $N^{\leqslant c_{\mathcal{B}^{\prime}}}_{G}(v)$, hence it depends only on the random values within $c_{\mathcal{B}^{\prime}}+1$ hops in $H$. Let $C_{1}=(c_{4}-c_{\mathcal{B}^{\prime}}-c_{2}-1)/(3c_{1})$, $C_{2}=c_{\mathcal{B}^{\prime}}+1$ and $C_{3}=10$; it is easy to verify that $C_{1}>C_{3}+4C_{2}+2$ for our choice of $c_{1},c_{2},c_{4},c_{\mathcal{B}^{\prime}}$. By Lemma 3.1, w.h.p., all the components of $H[N^{\leqslant c_{\mathcal{B}^{\prime}}}_{G}(\mathcal{A}^{\prime})]$ have size at most $O(\Delta(H)^{2c_{\mathcal{B}^{\prime}}+2}\log n)=O(\Delta^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}}\log n)$, which implies the claimed bound on the size of the largest connected component in $G_{\operatorname{dep}}$.
@@ 4.1 Graph Decomposition
4.2 The Color Coverage (CC) Property
In this section, we formally introduce the color coverage property, which we abbreviate CC, that was outlined in Section 2. Informally speaking we wish to ensure that the coloring outside of cliques $A_{i}$ looks random enough. Formally, we ask that each color is available to a constant fraction of the nodes on the inside of the cliques.
While $\mathrm{All}_{i}$ contains the nodes that are connected to all nodes in $A_{i}$, the nodes in $\operatorname{Big}_{i}^{+}$ are connected to most nodes (but still not too many) of $A_{i}$. Hence, coloring a single node in $\operatorname{Big}_{i}^{+}$ may bring $A_{i}$ close to not satisfying CC. But due to Lemma 4.1 the nodes in $\operatorname{Big}_{i}^{+}$ form a clique for $A_{i}$; hence, no two nodes in $\operatorname{Big}_{i}^{+}$ can receive the same color.
Property CC.
Consider some arbitrary step $j$ of a coloring algorithm. Let $\text{CC}_{j}(i,x)$ be the number of vertices in $A_{i}$ that have a neighbor $v\not\in A_{i}\cup\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ that gets colored with $x$ in step $j$. We say that CC is within budget for a single coloring step if $\text{CC}_{j}(i,x)<\Delta^{37/40}$ holds for every uncolored clique $A_{i}$ and for every color $x$. We say that CC holds for a clique $A_{i}$ at a point in time if $C$ has been within budget for every coloring step applied so far.
@@ 4.2 The Color Coverage (CC) Property
If we maintain CC throughout the coloring process for a clique $A_{i}$, then for all colors $x$ and at any point in the coloring process, at most $4\Delta/5$ vertices of $A_{i}$ have a neighbor outside of $A_{i}\cup\mathrm{All}_{i}$ with color $x$.
@@
Our algorithm has either $O(\log^{*}\Delta)$ or $O(\log\Delta)$ coloring steps [^6] In several of our coloring steps we compute a partial coloring by setting up an LLL that we solve via the shattering framework. That means that we color vertices in a pre-shattering and a post-shattering phase, both of which count as a separate coloring step in the context of this observation. In other words, the pre-shattering and post-shattering phases in these LLLs obtain a separate CC budget which simplifies the analysis significantly. , depending on the approach. Summing up the $\Delta^{37/40}$ budget of every individual step, results in $O(\log\Delta\cdot\Delta^{37/40})$ vertices in $A_{i}$ that have a $v\not\in A_{i}\cup\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ that is permanently colored with color $x$ (in any previous step). Consider a fixed color $x$. By Lemma 4.1-(c), $\operatorname{Big}_{i}^{+}$ forms a clique, and hence at most one node of $\operatorname{Big}_{i}^{+}$ can be colored $x$, and this vertex has at most $\tfrac{3}{4}\Delta+10^{8}\sqrt{\Delta}$ neighbors in $A_{i}$. Overall, the number of vertices $v\not\in A_{i}\cup\mathrm{All}_{i}$ adjacent to a vertex colored $x$ is bounded above by
@@
The high level idea of using Lemma 4.4 to show that CC is maintained for cliques in each step of our coloring procedure is as follows. Most of our coloring steps are based on some variant of random color trials, in which each node picks one (or multiple) candidate colors and then permanently retains one of its candidate colors if none of its neighbors chose it as a candidate. To prove that such a process maintains CC, fix a color $x\in[c]$, a clique $A_{i}$, and define a set $N_{v}$ for each vertex $v$ of the clique containing its external neighbors, except those in $\operatorname{Big}_{i}^{+}$. Let $Q$ be an upper bound on the sizes of the sets. As we exclude vertices in $\operatorname{Big}_{i}^{+}$ (those with more than $2\Delta^{9/10}$ neighbors in $A_{i}$) no vertex lies in more than $2\Delta^{9/10}$ sets. We consider a vertex $u\in\bigcup N_{v}$ marked if color $x$ is among its candidate colors. Clearly, the final colors picked by vertices are not independent, but the choices of candidate colors is. Hence, if the list of each vertex is of size $Q\cdot\Delta^{1/5}$ the properties of Lemma 4.4 are satisfied. We obtain that CC is maintained for color $x$ and clique $A_{i}$ with probability $\exp(-\Delta^{1/40})$.
@@ 4.3 Full Algorithm and Proofs of Theorems 1 and 2
Property (b) ensures that every vertex in $H$ has slack $U\cdot\Delta^{0.22}$, and importantly this slack remains as other vertices of the subgraph $H$ are colored. This allows us to color $H$ with $O(\log\Delta)$ iterated random color trial while maintaining Property CC. In Section 5 we prove the following lemma for coloring $\Pi$-ous subgraphs.
@@
Let $H$ be a $\Pi$-ous subgraph of $F-\bigcup_{i}A_{i}$. There is a $\mathsf{LOCAL}$ algorithm that, w.h.p., extends the coloring to all the vertices of $H$ in $\widetilde{O}(\log\Delta\cdot\log^{3}\log n)$ rounds while maintaining Property CC.
Note that when $\Delta=n^{\varepsilon}$, Lemma 4.5 does not guarantee a sublogarithmic runtime. For $\Delta\geqslant(\log n)^{50}$, we adapt the state-of-the-art $(\deg+1)$-list-coloring algorithm of [35] for coloring a $\Pi$-ous subgraph while maintaining CC; see Section 8 for details. Using Lemma 4.5 for low-degree graphs incurs an additional $O(\log\Delta)=O(\log\log n)$ factor in the runtime, but greatly simplifies the analysis.
@@
Let $H$ be a $\Pi$-ous subgraph of $F-\bigcup_{i}A_{i}$ and assume $\Delta\geqslant(\log n)^{50}$. There is a $\mathsf{LOCAL}$ algorithm that list-colors all vertices of $H$ in $O(\log^{*}n)$ rounds while maintaining Property CC.
We cannot use Lemma 4.5 directly to color the vertices of $S$ because they do not have enough initial slack to satisfy $\Pi$. As in [45], we begin by solving an LLL to provide them with $\Omega(\sqrt{\Delta})$ slack which is still not sufficient for $\Pi$. We then color the vertices of $S$ with Lemma 4.5 (or Lemma 4.6) in two batches to ensure that $\Pi$ holds, thereby that CC is maintained. Details can be found in Section 6.
@@
Let $c\geqslant\Delta-k_{\Delta}+1$ and $S$ be the set of vertices in $F$ as in Lemma 4.1-(1). There is a $\widetilde{O}(\log^{4}\log n)$-round distributed algorithm that, w.h.p., $c$-colors the vertices of $S$ while maintaining CC. When $\Delta\geqslant(\log n)^{50}$, it runs in $O(\log^{*}n)$ rounds.
As explained above, to color the cliques, we require that they satisfy CC. Since the cliques from $A_{L}$ are colored later, we must be careful about maintaining CC for those as we color $A_{H}$. We prove Lemma 4.8 in Section 7.
@@
Then, there is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., $c$-colors all the cliques in $A^{\prime}$ while maintaining CC for all the uncolored cliques outside $A^{\prime}$. For $\Delta\geqslant(\log n)^{50}$, it runs in $O(1)$ rounds.
@@
Coloring $A_{H}$. The collection of cliques in $A_{H}$ is colored via Lemma 4.8. We can apply the lemma as their CC was maintained by earlier coloring steps (Lemmas 4.7, 4.5 and 4.6). We emphasize that, by Lemma 4.8, CC of cliques in $A_{L}$ still holds after coloring cliques in $A_{H}$.
@@
Coloring $A_{L}$. The cliques in $A_{L}$ are also colored via Lemma 4.8, which can be done as CC holds for each of these cliques. Note that Lemma 4.8-(2) vacuously holds at this step because all the uncolored cliques are in $A^{\prime}=A_{L}$.

### 01 Apr 2026 v2

@@ 2 Technical Overview
2.1 Color Coverage ($C$) & $\Pi$-ous subgraphs
@@
To circumvent this issue, Molloy and Reed constrain all earlier steps of the coloring algorithm so that the coloring outside each clique behaves sufficiently randomly. Concretely, this means that for every clique and every color, only a small number of vertices are forbidden from using that color. We formalize this requirement as the color coverage (henceforth $C$) property. Ensuring that $C$ holds until coloring the dense parts of the graph is essential to completing the coloring. See Section 4.2 for the precise definition of $C$.
@@
The $C$ condition can be maintained if all earlier coloring steps are sufficiently random. If each dense node has at most $U$ neighbors outside its clique, and each of them picks one random color out of $U\cdot\Delta^{0.22}$ colors, then few of these outside neighbors pick the same color, thereby maintaining the $C$ condition. Thus, the key to ensuring the $C$ condition in each step of the algorithm is to maintain large lists of available colors $L(v)$ (colors that do not currently appear in the neighborhood of $v$) compared to the number of outside neighbors of the cliques. We capture this condition through $\Pi$, which we state in a simplified form as follows:
@@
Subgraphs satisfying $\Pi$ are called $\Pi$-ous (pronounced pious). We can color a $\Pi$-ous subgraph $H$ while maintaining $C$ by a combination of iterated random color trials (Rct) and multi color trials (Mct). Condition (b) of $\Pi$ ensures that the list of available colors for each vertex of $H$ has $U\cdot\Delta^{0.22}$ slack that remains and ensures a sufficiently large list regardless of how the coloring of $H$ proceeds. Slack does not decrease because coloring a neighbor removes both one competing vertex and one available color, leaving their difference—the slack—unchanged or increased.
@@ 2.2 Overview of the Coloring Algorithm
Our algorithm centers on identifying $\Pi$-ous subgraphs that can be colored while maintaining $C$, thereby ensuring that the dense vertices can be colored last. We build on the structural decomposition of Molloy and Reed, which nearly [^4] The subgraphs induced by $B_{H}$ and $B_{L}$ are $\Pi$-ous by construction, however the subgraph induced by $S$ is not $\Pi$-ous right away; see Section 6 for more details. partitions the sparser regions of the graph into $\Pi$-ous subgraphs [45]. More precisely, they reduced the problem of finding a $c$-coloring of $G$ to the problem of finding a $c$-coloring of a graph $F$ with a very specific structure. Bamas and Esperet also showed that $F$ can be constructed from $G$ in $O(1)$ rounds in the $\mathsf{LOCAL}$ model, and that one can efficiently recover a $c$-coloring of $G$ from a $c$-coloring of $F$ in the distributed setting [4].
@@
To achieve a sublogarithmic distributed runtime, we color cliques in $A_{H}$ and $A_{L}$ respectively in parallel. For each clique $A_{i}$, we assign a random permutation of the colors to its vertices avoiding any conflicts inside the clique. With this process, only $O(\sqrt{\Delta})$ vertices per clique conflict with external neighbors. Each such unhappy vertex $v$ seeks a partner $u$ in $A_{i}$ so that both become happy if they swap their colors. $C$ crucially guarantees a large supply of valid swap candidates.
@@ 2.3 Smaller Degrees and the Lovász Local Lemma
As one randomly colors vertices in the pre-shattering phase, one may be forced to uncolor certain vertices. For instance, if a color appears too frequently around a clique and $C$ is violated. This may interfere with the progress measures of our algorithm, which are typically formulated as additional bad events. For instance, uncoloring vertices may violate an event ensuring a constant-factor decrease in the uncolored degree. Such events are usually resolved during the post-shattering phase.
@@
Since we only pick candidates in pre- and post-shattering using separate budgets, this increases the number of neighbors with a given color by a factor of at most two compared to having one single budget for both steps. Conveniently, the same analysis bounds the probability of breaking $C$ for the pre- and post-shattering.
@@ 3. III:
By (S2), (S4), and the union bound over all sets AA with $v$ as a vertex, we have that
@@
By (S3), whether $A\in\mathcal{A}^{\prime}$ depend only on the random variables $X_{u}$ of vertices in AA, which has diameter one in $H$. Whether $v\in N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}}(\mathcal{A}^{\prime})$ depends on whether $A\in\mathcal{A}^{\prime}$ for all the AA with a vertex in $N^{\leqslant c_{\mathcal{B}^{\prime}}}_{G}(v)$, hence it depends only on the random values within $c_{\mathcal{B}^{\prime}}+1$ hops in $H$. Let $C_{1}=(c_{4}-c_{\mathcal{B}^{\prime}}-c_{2}-1)/(3c_{1})$, $C_{2}=c_{\mathcal{B}^{\prime}}+1$ and $C_{3}=10$; it is easy to verify that $C_{1}>C_{3}+4C_{2}+2$ for our choice of $c_{1},c_{2},c_{4},c_{\mathcal{B}^{\prime}}$. By Lemma 3.1, w.h.p., all the components of $H[N^{\leqslant c_{\mathcal{B}^{\prime}}}_{G}(\mathcal{A}^{\prime})]$ have size at most $O(\Delta(H)^{2c_{\mathcal{B}^{\prime}}+2}\log n)=O(\Delta^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}}\log n)$, which implies the claimed bound on the size of the largest connected component in $G_{\operatorname{dep}}$.
@@ 4.1 Graph Decomposition
4.2 The Color Coverage ($C$) Property
In this section, we formally introduce the color coverage property, which we abbreviate $C$, that was outlined in Section 2. Informally speaking we wish to ensure that the coloring outside of cliques $A_{i}$ looks random enough. Formally, we ask that each color is available to a constant fraction of the nodes on the inside of the cliques.
While $\mathrm{All}_{i}$ contains the nodes that are connected to all nodes in $A_{i}$, the nodes in $\operatorname{Big}_{i}^{+}$ are connected to most nodes (but still not too many) of $A_{i}$. Hence, coloring a single node in $\operatorname{Big}_{i}^{+}$ may bring $A_{i}$ close to not satisfying $C$. But due to Lemma 4.1 the nodes in $\operatorname{Big}_{i}^{+}$ form a clique for $A_{i}$; hence, no two nodes in $\operatorname{Big}_{i}^{+}$ can receive the same color.
Property $C$.
Consider some arbitrary step $j$ of a coloring algorithm. Let $\text{CC}_{j}(i,x)$ be the number of vertices in $A_{i}$ that have a neighbor $v\not\in A_{i}\cup\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ that gets colored with $x$ in step $j$. We say that $C$ is within budget for a single coloring step if $\text{CC}_{j}(i,x)<\Delta^{37/40}$ holds for every uncolored clique $A_{i}$ and for every color $x$. We say that $C$ holds for a clique $A_{i}$ at a point in time if $C$ has been within budget for every coloring step applied so far.
@@ 4.2 The Color Coverage ($C$) Property
If we maintain $C$ throughout the coloring process for a clique $A_{i}$, then for all colors $x$ and at any point in the coloring process, at most $4\Delta/5$ vertices of $A_{i}$ have a neighbor outside of $A_{i}\cup\mathrm{All}_{i}$ with color $x$.
@@
Our algorithm has either $O(\log^{*}\Delta)$ or $O(\log\Delta)$ coloring steps [^6] In several of our coloring steps we compute a partial coloring by setting up an LLL that we solve via the shattering framework. That means that we color vertices in a pre-shattering and a post-shattering phase, both of which count as a separate coloring step in the context of this observation. In other words, the pre-shattering and post-shattering phases in these LLLs obtain a separate $C$ budget which simplifies the analysis significantly. , depending on the approach. Summing up the $\Delta^{37/40}$ budget of every individual step, results in $O(\log\Delta\cdot\Delta^{37/40})$ vertices in $A_{i}$ that have a $v\not\in A_{i}\cup\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ that is permanently colored with color $x$ (in any previous step). Consider a fixed color $x$. By Lemma 4.1-(c), $\operatorname{Big}_{i}^{+}$ forms a clique, and hence at most one node of $\operatorname{Big}_{i}^{+}$ can be colored $x$, and this vertex has at most $\tfrac{3}{4}\Delta+10^{8}\sqrt{\Delta}$ neighbors in $A_{i}$. Overall, the number of vertices $v\not\in A_{i}\cup\mathrm{All}_{i}$ adjacent to a vertex colored $x$ is bounded above by
@@
The high level idea of using Lemma 4.4 to show that $C$ is maintained for cliques in each step of our coloring procedure is as follows. Most of our coloring steps are based on some variant of random color trials, in which each node picks one (or multiple) candidate colors and then permanently retains one of its candidate colors if none of its neighbors chose it as a candidate. To prove that such a process maintains $C$, fix a color $x\in[c]$, a clique $A_{i}$, and define a set $N_{v}$ for each vertex $v$ of the clique containing its external neighbors, except those in $\operatorname{Big}_{i}^{+}$. Let $Q$ be an upper bound on the sizes of the sets. As we exclude vertices in $\operatorname{Big}_{i}^{+}$ (those with more than $2\Delta^{9/10}$ neighbors in $A_{i}$) no vertex lies in more than $2\Delta^{9/10}$ sets. We consider a vertex $u\in\bigcup N_{v}$ marked if color $x$ is among its candidate colors. Clearly, the final colors picked by vertices are not independent, but the choices of candidate colors is. Hence, if the list of each vertex is of size $Q\cdot\Delta^{1/5}$ the properties of Lemma 4.4 are satisfied. We obtain that $C$ is maintained for color $x$ and clique $A_{i}$ with probability $\exp(-\Delta^{1/40})$.
@@ 4.3 Full Algorithm and Proofs of Theorems 1 and 2
Property (b) ensures that every vertex in $H$ has slack $U\cdot\Delta^{0.22}$, and importantly this slack remains as other vertices of the subgraph $H$ are colored. This allows us to color $H$ with $O(\log\Delta)$ iterated random color trial while maintaining Property $C$. In Section 5 we prove the following lemma for coloring $\Pi$-ous subgraphs.
@@
Let $H$ be a $\Pi$-ous subgraph of $F-\bigcup_{i}A_{i}$. There is a $\mathsf{LOCAL}$ algorithm that, w.h.p., extends the coloring to all the vertices of $H$ in $\widetilde{O}(\log\Delta\cdot\log^{3}\log n)$ rounds while maintaining Property $C$.
Note that when $\Delta=n^{\varepsilon}$, Lemma 4.5 does not guarantee a sublogarithmic runtime. For $\Delta\geqslant(\log n)^{50}$, we adapt the state-of-the-art $(\deg+1)$-list-coloring algorithm of [35] for coloring a $\Pi$-ous subgraph while maintaining $C$; see Section 8 for details. Using Lemma 4.5 for low-degree graphs incurs an additional $O(\log\Delta)=O(\log\log n)$ factor in the runtime, but greatly simplifies the analysis.
@@
Let $H$ be a $\Pi$-ous subgraph of $F-\bigcup_{i}A_{i}$ and assume $\Delta\geqslant(\log n)^{50}$. There is a $\mathsf{LOCAL}$ algorithm that list-colors all vertices of $H$ in $O(\log^{*}n)$ rounds while maintaining Property $C$.
We cannot use Lemma 4.5 directly to color the vertices of $S$ because they do not have enough initial slack to satisfy $\Pi$. As in [45], we begin by solving an LLL to provide them with $\Omega(\sqrt{\Delta})$ slack which is still not sufficient for $\Pi$. We then color the vertices of $S$ with Lemma 4.5 (or Lemma 4.6) in two batches to ensure that $\Pi$ holds, thereby that $C$ is maintained. Details can be found in Section 6.
@@
Let $c\geqslant\Delta-k_{\Delta}+1$ and $S$ be the set of vertices in $F$ as in Lemma 4.1-(1). There is a $\widetilde{O}(\log^{4}\log n)$-round distributed algorithm that, w.h.p., $c$-colors the vertices of $S$ while maintaining $C$. When $\Delta\geqslant(\log n)^{50}$, it runs in $O(\log^{*}n)$ rounds.
As explained above, to color the cliques, we require that they satisfy $C$. Since the cliques from $A_{L}$ are colored later, we must be careful about maintaining $C$ for those as we color $A_{H}$. We prove Lemma 4.8 in Section 7.
@@
Then, there is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., $c$-colors all the cliques in $A^{\prime}$ while maintaining $C$ for all the uncolored cliques outside $A^{\prime}$. For $\Delta\geqslant(\log n)^{50}$, it runs in $O(1)$ rounds.
@@
Coloring $A_{H}$. The collection of cliques in $A_{H}$ is colored via Lemma 4.8. We can apply the lemma as their $C$ was maintained by earlier coloring steps (Lemmas 4.7, 4.5 and 4.6). We emphasize that, by Lemma 4.8, $C$ of cliques in $A_{L}$ still holds after coloring cliques in $A_{H}$.
@@
Coloring $A_{L}$. The cliques in $A_{L}$ are also colored via Lemma 4.8, which can be done as $C$ holds for each of these cliques. Note that Lemma 4.8-(2) vacuously holds at this step because all the uncolored cliques are in $A^{\prime}=A_{L}$.